OpenStudy (anonymous):

17x^2-10x-15=0 List all the possible rational zeros

OpenStudy (anonymous):

add fractions \(\frac{p}{q}\) where p divides 15 and q divides 17 there are a bunch \[\{\pm1,\pm3,\pm5,\pm15,\pm\frac{1}{17},\pm\frac{5}{17},\pm\frac{15}{17}\}\]